# The Fourier spectral approach to the spatial discretization of quasilinear hyperbolic systems

### v1

@@ 3.1 Discretization with smooth low-pass filters
We consider the spatial discretization with smooth low-pass filters first, as it is more similar to the previous sections. Recall that the spatial discretization with the smooth low-pass filter ${\sf S}_{N}$ amounts to solving
@@
where $A_{j}^{0}\coloneq A_{j}(\bm{0})$ and $A_{j}^{1}(\bm{U})\coloneq A_{j}(\bm{U})-A_{j}^{0}$. However, we face a difficulty that while any symmetrizer for the system (1.1), $S(\bm{U})$, readily provides a suitable symmetrizer of the semi-discretized system (3.1), such is not the case for system (3.2), and additional assumptions are needed.
@@
Supposing the Assumptions A.1 and A.2 hold, and decomposing $A_{j}(\bm{U})=A_{j}^{0}+A_{j}^{1}(\bm{U})$ where $A_{j}^{0}\coloneq A_{j}(\bm{0})$ and $S(\bm{U})=S^{0}+S^{1}(\bm{U})$ where $S^{0}\coloneq S(\bm{0})$, we have
$\forall j\in\{1,\ldots,d\},\ \forall\bm{U}\in{\mathcal{U}},\quad S^{0}A_{j}^{0},\ S^{0}A_{j}^{1}(\bm{U})+S^{1}(\bm{U})A_{j}^{0}\text{ and }S^{1}(\bm{U})A_{j}^{1}(\bm{U})\text{ are symmetric.}$.
@@
As can be seen by Taylor-expanding $S(\bm{U})A_{j}(\bm{U})$ about the origin, Assumption A.3 holds in particular when $A_{j}(\bm{U})$ and $S(\bm{U})$ are linear, that is $A_{j}^{1}(\bm{U})$ and $S^{1}(\bm{U})$ are homogeneous polynomials of order $1$ in the entries of $\bm{U}$; see also Remark 3.5.
@@
Suppose that the system (1.1) satisfies Assumptions A.1 and A.2. Let ${s>d/2+1}$, $M>0$ and ${\mathcal{K}}\subset{\mathcal{U}}$ compact. There exists $C>0$ and $T>0$ (depending only on $s,M,$ and ${\mathcal{K}}$) such that for every $N\in\mathbb{N}$ and for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}\leq M$ and taking values in ${\mathcal{K}}$, there exists a unique $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ maximal-in-time classical solution to (3.1) and $\bm{U}\big{|}_{t=0}={\sf P}_{N}\bm{U}^{0}$. The open time interval $I_{N}\supset[0,T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}]$ and for all $0\leq t\leq T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}$,
@@
$\left|({\rm Sym}(\bm{U}_{N})\partial_{x_{j}}\dot{\bm{U}}_{N},\dot{\bm{U}}_{N})_{L^{2}}\right|\leq C({\mathcal{K}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{U}_{N}\right|_{H^{s^{\prime}}}^{2}.$
@@
$\left|({\rm Com}(\bm{U}_{N})\partial_{x_{j}}\dot{\bm{U}}_{N},\dot{\bm{U}}_{N})_{L^{2}}\right|\leq C({\mathcal{K}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{U}_{N}\right|_{H^{s^{\prime}}}^{2}.$
@@
The first two terms can be estimated in the standard manner using that $S(\bm{U})$ is a symmetrizer, i.e. Assumption A.2. The third term is estimated using the regularizing properties of the commutator with ${\sf S}_{N/2}^{1/2}$, Proposition A.6. The fourth term is estimated thanks to Assumption A.1. Altogether we obtain
@@
With this estimate in hand, we can follow the proof of Proposition 2.4. Using that $\|\operatorname{Id}-{\sf S}_{N/2}^{1/2}\|_{H^{s}\to L^{2}}\leq\langle N/4\rangle^{-s}$, the coercivity of $S(\bm{U}_{N})$ —see (3.4)— and the uniform estimates for $\bm{U}_{N}$ stated in Proposition 3.3 we infer (by Grönwall’s lemma), denoting $M^{*}\coloneq 2\sup_{t\in I^{*}}\left|\bm{U}\right|_{H^{s^{\prime}}}$ and assuming $\sup_{t\in I^{*}}\left|\bm{U}_{N}(t,\cdot)\right|_{H^{s}}\leq M^{*}$ and $\bm{U}_{N}(I^{\star}\times\mathbb{R}^{n})\subset{\mathcal{K}}^{*}\subset{\mathcal{U}}$,
@@
This yields the desired estimate for $r=0$, and the general case follows by interpolation. A continuity argument concludes the proof as in Proposition 2.4. ∎
@@
Let us comment on the restrictive Assumption A.3 arising in the study of system (3.2). This assumption allows to ensure that we can construct a symmetrizer operator $\tilde{S}(\bm{U})$ that is
@@
    such that $\tilde{S}(\bm{U})(A_{j}^{0}+{\sf S}_{N}A_{j}^{1}(\bm{U}))$ is self-adjoint up to regularizing operators of order $-1$.
@@
$\tilde{S}(\bm{U})[\circ]\coloneq\sum_{\ell=0}^{m}{\sf S}_{N}^{\ell/2}(S^{\ell}(\bm{U}){\sf S}_{N}^{\ell/2}[\circ]).$
@@
where entries of $S^{\ell}(\cdot)$ and $A_{j}^{\ell}(\cdot)$ are homogeneous polynomials of order $\ell$.
By Taylor expansion about $\bm{U}=\bm{0}$ of the symmetric matrices $S(\bm{U})$ and $S(\bm{U})A_{j}(\bm{U})$ we find that $S^{\ell}(\bm{U})$ and $\sum_{\ell_{1}+\ell_{2}=\ell}S^{\ell_{1}}(\bm{U})A_{j}^{\ell_{2}}(\bm{U})$ are symmetric. This shows that $\tilde{S}(\bm{U})$ is self-adjoint for the $L^{2}(\mathbb{R}^{d})^{n}$ inner-product and, denoting $\tilde{A}_{j}(\bm{U})[\circ]\coloneq\sum_{\ell=0}^{m_{j}}{\sf S}_{N}^{\ell}(A_{j}^{\ell}(\bm{U}_{N})[\circ])$, $\tilde{S}(\bm{U})\tilde{A}_{j}(\bm{U})$ is self-adjoint up to regularizing operators of order $-1$. Notice however that some additional restrictions on smallness of $\bm{U}\in{\mathcal{U}}$ may be necessary to enforce the coercivity of the operator $\tilde{S}(\bm{U})$.
@@ 3.2 Discretization with sharp low-pass filters
The analysis of the previous section fails, due to the lack of good commutator properties of the operator ${\sf P}_{N}$. Specifically, when applying a symmetrizer $S(\bm{U})$ to the underlying system (1.1) we may write
@@
Then we remark that (1.1) takes the Hamiltonian form
@@
$-\sum_{j=1}^{d}\Big{(}S_{j}^{0}{\sf P}_{N}\big{(}[\partial_{x_{j}},S(\bm{U}_{N})]\Lambda^{s^{\prime}}\bm{U}_{N}\big{)},{\sf P}_{N}\big{(}S(\bm{U}_{N})\Lambda^{s^{\prime}}\bm{U}_{N}\big{)}\Big{)}_{L^{2}}$
@@ 4 Numerical experiments for the Saint-Venant system
which describes the propagation of shallow water waves in the flat-bottom situation; see [17]. Specifically, the scalar variable $\eta$ describes the elevation of the surface of a layer of homogeneous, incompressible and inviscid fluid and the variable $\bm{u}$ represents the layer-averaged horizontal velocity of fluid particles (both depending on time and horizontal space). The gravitational constant has been set to $g=1$.
@@
Thanks to these features, numerical experimenyts on the Saint-Venant systems allow to showcase our numerical findings, that we summarize here for the sake of readability.
@@ 4.1 Analysis of the Saint-Venant system
We can apply the analysis of the previous sections to systems (4.1) and (4.2) due to the following result.
@@
[…]
@@
[…]
$S_{{\mathscr{H}}}((\eta,\bm{u}))=\begin{pmatrix}1&\bm{u}^{\top}\\ \bm{u}&(1+\eta)\operatorname{Id}\end{pmatrix}\text{ and }S_{j}^{0}=\begin{pmatrix}0&\bm{e}_{j}^{\top}\\ \bm{e}_{j}&\bm{0}\end{pmatrix}\text{ where }\bm{e}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},\bm{e}_{2}=\begin{pmatrix}0\\ 1\end{pmatrix},$ (4.4)
@@
[…]
@@
As discussed in the previous section, Section 3.2, we require more stringent structural assumptions to show convergence for symmetrizable systems when discretizing with the sharp low-pass filter ${\sf P}_{N}$. The Saint-Venant system when $d=1$ illustrates that we may also have to require a stricter condition on the inital data ($\bm{U}^{0}$ taking values in ${\mathcal{U}}_{{\mathscr{H}}}$), even for systems that satify the structural assumptions.
This is because any symmetrizer for the underlying system in the sense of Assumption A.2 satisfying the additional compatibility Assumption A.3 can be used to construct a symmetrizer to the semi-discrete systems (4.6) and (4.7). For the the semi-discretization with ${\sf P}_{N}$ (4.8) on the other hand, we use the symmetrizer directly related to the structure of the system through Assumption A.5 to construct a symmetrizer for (4.8).
@@
Let $\bm{U}\coloneq(\eta_{p},u_{p})$ where $\eta_{p}\coloneq-\frac{1}{2}\cos(px)$ and $u_{p}\coloneq\sin(px)$, and $\bm{V}_{N}\coloneq(0,\sin((N-q)x))$ with $0\leq q<p\ll N$. A direct calculation yields
$(\operatorname{Id}-{\sf P}_{N})\big{(}A(\bm{U})(\partial_{x}{\sf P}_{N}\bm{V}_{N})\big{)}$ $=\frac{N-q}{4}\begin{pmatrix}-\cos((N-q+p)x)\\ 2\sin((N-q+p)x)\end{pmatrix}+{\mathcal{O}}(1),$
@@
${\sf P}_{N}\Big{(}S(\bm{U})(\operatorname{Id}-{\sf P}_{N})\big{(}A(\bm{U})(\partial_{x_{j}}{\sf P}_{N}\bm{V}_{N})\big{)}\Big{)}=\frac{N-q}{8}\begin{pmatrix}0\\ -\sin((N-q)x)\end{pmatrix}+{\mathcal{O}}(1),$
@@
$J_{N}(\bm{U},\bm{V}_{N})=-\frac{\pi}{8}(N-q)+{\mathcal{O}}(1).$
This shows that one cannot propagate for positive time (at least in a direct manner) a uniform-in-$N$ control of the energy functional ${\mathcal{F}}_{s}(\bm{U}_{N})$ for $\bm{U}_{N}$ the solution emerging from initial data $\bm{U}_{N}^{0}\coloneq\bm{U}+\bm{V}/(N-q)^{s}$, despite the fact that $\bm{U}_{N}^{0}\in{\mathcal{U}}$ since $01+\eta_{N}\geq 1/2>$ 0, and $\left|\bm{U}_{N}^{0}\right|_{H^{s}}\approx 1$. Notice also that $\bm{U}_{N}^{0}\notin{\mathcal{U}}_{{\mathscr{H}}}$ since $01+\eta(\frac{\pi}{2p})-|u(\frac{\pi}{2p})|^{2}=$ 0, but one could enforce $\bm{U}_{N}^{0}\in{\mathcal{U}}_{{\mathscr{H}}}$ (while keeping valid all previous statements) by considering for instance $u_{p}=\frac{1}{2}\sin(px)$. In that case, the Hamiltonian structure allows to propagate the functional ${\mathcal{F}}_{{\mathscr{H}},s}(\bm{U})\approx\big{\lvert}\bm{U}_{N}\big{\rvert}_{H^{s}}$ with

### v2

@@ 3.1 Discretization with smooth low-pass filters
We consider the spatial discretization with smooth low-pass filters first, as it is more similar to the previous section. Recall that the spatial discretization with the smooth low-pass filter ${\sf S}_{N}$ amounts to solving
@@
where $A_{j}^{0}\coloneq A_{j}(\bm{0})$ and $A_{j}^{1}(\cdot)\coloneq A_{j}(\cdot)-A_{j}^{0}$. However, we face a difficulty that while any symmetrizer for the system (1.1), $S(\cdot)$, readily provides a suitable symmetrizer of the semi-discretized system (3.1), such is not the case for system (3.2), and additional assumptions are needed.
@@
Supposing the Assumptions A.1 and A.2 hold, and decomposing $A_{j}(\cdot)=A_{j}^{0}+A_{j}^{1}(\cdot)$ where $A_{j}^{0}\coloneq A_{j}(\bm{0})$ and $S(\cdot)=S^{0}+S^{1}(\cdot)$ where $S^{0}\coloneq S(\bm{0})$, we have
$\forall j\in\{1,\ldots,d\},\ \forall\bm{U}\in{\mathcal{U}},\quad S^{0}A_{j}^{0},\ S^{0}A_{j}^{1}(\bm{U})+S^{1}(\bm{U})A_{j}^{0}\text{ and }S^{1}(\bm{U})A_{j}^{1}(\bm{U})\text{ are symmetric.}$
@@
As can be seen by Taylor-expanding $S(\bm{U})A_{j}(\bm{U})$ about the origin, Assumption A.3 holds in particular when $A_{j}(\cdot)$ and $S(\cdot)$ are linear, that is entries of $A_{j}^{1}(\cdot)$ and $S^{1}(\cdot)$ are homogeneous polynomials of degree 1; see also Remark 3.6.
@@
Suppose that the system (1.1) satisfies Assumptions A.1 and A.2. Let s>d/2+1{s>d/2+1}, $M>0$ and ${\mathcal{K}}\subset{\mathcal{U}}$ compact. There exists $C>0$ and $T>0$ (depending only on $s,M,$ and $\mathcal{K}$}) such that for every $N\in\mathbb{N}$ and for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}\leq M$ and taking values in $\mathcal{K}$}, there exists a unique $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ maximal-in-time classical solution to (3.1) and $\bm{U}\big|_{t=0}={\sf P}_{N}\bm{U}^{0}$. The open time interval $I_{N}\supset[0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}]$ and for all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$,
@@
$\left|\big({\rm Sym}(\bm{U}_{N})\partial_{x_{j}}\dot{\bm{U}}_{N},\dot{\bm{U}}_{N}\big)_{L^{2}}\right|\leq C({\mathcal{K}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{U}_{N}\right|_{H^{s^{\prime}}}^{2}.$
@@
$\left|\big({\rm Com}(\bm{U}_{N})\partial_{x_{j}}\dot{\bm{U}}_{N},\dot{\bm{U}}_{N}\big)_{L^{2}}\right|\leq C({\mathcal{K}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{U}_{N}\right|_{H^{s^{\prime}}}^{2}.$
@@
The first two terms can be estimated in the standard manner using that $S(\cdot)$ is a symmetrizer, i.e. Assumption A.2. The third term is estimated using the regularizing properties of the commutator with ${\sf S}_{N/2}^{1/2}$, Proposition A.6. The fourth term is estimated thanks to Assumption A.1. Altogether we obtain
@@
With this estimate in hand, we can follow the proof of Proposition 2.4. Using that $\|\operatorname{Id}-{\sf S}_{N/2}^{1/2}\|_{H^{s}\to L^{2}}\leq\langle N/4\rangle^{-s}$, the coercivity of $S(\bm{U}_{N})$ —see (3.4)— and the uniform estimates for $\bm{U}_{N}$ stated in Proposition 3.4 we infer (by Grönwall’s lemma), denoting $M^{*}\coloneq 2\sup_{t\in I^{*}}\left|\bm{U}(t,\cdot)\right|_{H^{s_{0}}}$ with $d/2+1<s_{0}<s$ and assuming $\sup_{t\in I^{*}}\left|\bm{U}_{N}(t,\cdot)\right|_{H^{s_{0}}}\leq M^{*}$ and $\bm{U}_{N}(I^{*}\times\mathbb{R}^{n})\subset{\mathcal{K}}^{*}\subset{\mathcal{U}}$,
@@
This yields the desired estimate for $r=0$, and the general case follows by interpolation. A continuity argument as in Proposition 2.4 allows to secure the bound $\sup_{t\in I^{*}}\left|\bm{U}_{N}(t,\cdot)\right|_{H^{s_{0}}}\leq M^{*}$ and the assumption $\bm{U}_{N}(I^{*}\times\mathbb{R}^{n})\subset{\mathcal{K}}^{*}\subset{\mathcal{U}}$} for $N$ sufficiently large (by the convergence $\bm{U}_{N}(t,\cdot)\to\bm{U}(t,\cdot)$ as $N\to\infty$ in $H^{s_{0}}(\mathbb{R}^{d})^{n}\subset W^{1,\infty}(\mathbb{R}^{d})^{n}$) along the desired estimate, which concludes the proof . ∎
@@
[…]
@@
    such that $\tilde{S}(\bm{U})(A_{j}^{0}+{\sf S}_{N}A_{j}^{1}(\bm{U}))$ is self-adjoint up to regularizing operators of order 1-1.
@@
$\tilde{S}(\bm{U})[\circ]\coloneq\sum_{\ell=0}^{m}{\sf S}_{N}^{\ell/2}(S^{\ell}(\bm{U}){\sf S}_{N}^{\ell/2}[\circ]),$
@@
where entries of $S^{\ell}(\cdot)$ and $A_{j}^{\ell}(\cdot)$ are homogeneous polynomials of degree $\ell$.
By Taylor expansion about the origin of $S(\cdot)$ and $S(\cdot)A_{j}(\cdot)$ and homogeneity we find that for all $\bm{U}\in{\mathcal{U}}$, $S^{\ell}(\bm{U})$ and $\sum_{\ell_{1}+\ell_{2}=\ell}S^{\ell_{1}}(\bm{U})A_{j}^{\ell_{2}}(\bm{U})$ are symmetric. This shows that $\tilde{S}(\bm{U})$ is self-adjoint for the $L^{2}(\mathbb{R}^{d})^{n}$ inner-product and, denoting $\tilde{A}_{j}(\bm{U})[\circ]\coloneq\sum_{\ell=0}^{m_{j}}{\sf S}_{N}^{\ell}(A_{j}^{\ell}(\bm{U})[\circ])$, one has that $\tilde{S}(\bm{U})\tilde{A}_{j}(\bm{U})$ is self-adjoint up to regularizing operators of order 1-1. Notice however that some additional restrictions on $\bm{U}\in{\mathcal{U}}$ may be necessary to enforce the coercivity of the operator $\tilde{S}(\bm{U})$.
@@ 3.2 Discretization with sharp low-pass filters
The analysis of the previous section fails, due to the lack of good commutator properties of the operator ${\sf P}_{N}$. Specifically, when applying a symmetrizer $S(\cdot)$ to the underlying system (1.1) we may write
@@
Then we remark that under the Assumption A.5 (1.1) takes the Hamiltonian form
@@
$+\sum_{j=1}^{d}\Big(S_{j}^{0}{\sf P}_{N}\big([\partial_{x_{j}},S(\bm{U}_{N})]\Lambda^{s^{\prime}}\bm{U}_{N}\big),{\sf P}_{N}\big(S(\bm{U}_{N})\Lambda^{s^{\prime}}\bm{U}_{N}\big)\Big)_{L^{2}}$
@@ 4 Numerical experiments for the Saint-Venant system
which describes the propagation of shallow water waves in the flat-bottom situation; see [17]. Specifically, the scalar variable $\eta$ describes the elevation of the surface of a layer of homogeneous, incompressible and inviscid fluid and the variable $\bm{u}$ represents the layer-averaged horizontal velocity of fluid particles (both depending on time and horizontal space). The gravitational constant and reference depth have been set to $g=1$ and $H=1$.
@@
Thanks to these features, numerical experiments on the Saint-Venant systems allow to showcase our numerical findings, that we summarize here for the sake of readability.
@@ 4.1 Analysis of the Saint-Venant system
We can apply the analysis of the previous section to systems (4.1) and (4.2) due to the following result.
@@
System (4.1) is a symmetrizable hyperbolic system in the sense of Assumptions A.1 and A.2 with hyperbolic domain ${\mathcal{U}}\coloneq\{(\eta,\bm{u})\in\mathbb{R}^{1+d}\ :\ 1+\eta>0\}$ and
@@
System (4.2) satisfies Assumption A.5 (and hence Assumptions A.1 and A.2; see Remark 3.7) with hyperbolic domain ${\mathcal{U}}_{{\mathscr{H}}}\coloneq\{(\eta,\bm{u})\in\mathbb{R}^{1+d}\ :\ 1+\eta-\left|\bm{u}\right|^{2}>0\}$ and
$S_{{\mathscr{H}}}((\eta,\bm{u}))=\begin{pmatrix}1&\bm{u}^{\top}\\ \bm{u}&(1+\eta)\operatorname{Id}\end{pmatrix}\text{ and }S_{j}^{0}=\begin{pmatrix}0&\bm{e}_{j}^{\top}\\ \bm{e}_{j}&\bm{0}\end{pmatrix}\text{ where }\bm{e}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},\bm{e}_{2}=\begin{pmatrix}0\\ 1\end{pmatrix}$ (4.4)
@@
(Convergence) The statement of Proposition 3.5, concerning spectral convergence of solutions to the semi-discrete systems (4.6) and (4.7), holds for the Saint-Venant system (4.1) whenever $\bm{U}^{0}$ takes values in ${\mathcal{U}}\coloneq\{(\eta,\bm{u})\in\mathbb{R}^{1+d}\ :\ 1+\eta>0\}$}. If additionally $\bm{U}^{0}$ takes values in ${\mathcal{U}}_{{\mathscr{H}}}\coloneq\{(\eta,\bm{u})\in\mathbb{R}^{1+d}\ :\ 1+\eta-\left|\bm{u}\right|^{2}>0\}$, then Proposition 3.5 also holds for the Hamiltonian Saint-Venant system (4.2).
@@
As discussed in the previous section, Section 3.2, we require more stringent structural assumptions to show convergence for symmetrizable systems when discretizing with the sharp low-pass filter ${\sf P}_{N}$. The Saint-Venant system when $d=1$ illustrates that we may also have to impose more stringent restrictions on the inital data (namely $\bm{U}^{0}$ taking values in $\mathcal{U}$$_{{\mathscr{H}}}$), even for systems that satify the structural assumptions.
This is because any symmetrizer for the underlying system in the sense of Assumption A.2 satisfying the additional compatibility Assumption A.3 can be used to construct a symmetrizer to the semi-discrete systems (4.6) and (4.7). For the semi-discretization with ${\sf P}_{N}$ (4.8) on the other hand, we use the symmetrizer directly related to the structure of the system through Assumption A.5.
@@
Let $\bm{U}\coloneq(\eta_{p},u_{p})$ where $\eta_{p}(x)\coloneq-\frac{1}{2}\cos(px)$, $u_{p}(x)\coloneq\sin(px)$ and $\bm{V}_{N}\coloneq(0,v_{N,q})$ where $v_{N,q}(x)\coloneqq\sin((N-q)x)$} with $0\leq q<p\ll N$. A direct calculation yields
$(\operatorname{Id}-{\sf P}_{N})\big(A(\bm{U})(\partial_{x}{\sf P}_{N}\bm{V}_{N})\big)$ $=\frac{N-q}{4}\begin{pmatrix}-\cos((N-q+p)x)\\ 2\sin((N-q+p)x)\end{pmatrix}$,
@@
${\sf P}_{N}\Big(S(\bm{U})(\operatorname{Id}-{\sf P}_{N})\big(A(\bm{U})(\partial_{x_{j}}{\sf P}_{N}\bm{V}_{N})\big)\Big)=\frac{N-q}{8}\begin{pmatrix}0\\ -\sin((N-q)x)\end{pmatrix}$,
@@
$J_{N}(\bm{U},\bm{V}_{N})=-\frac{\pi}{8}(N-q).$
This shows that one cannot propagate for positive time (at least in a direct manner) a uniform-in-$N$ control of the energy functional ${\mathcal{F}}_{s}(\bm{U}_{N})$ for $\bm{U}_{N}$ the solution emerging from initial data $\bm{U}_{N}^{0}\coloneq\bm{U}+\bm{V}_{N}/(N-q)^{s}$, despite the fact that $\bm{U}_{N}^{0}\in{\mathcal{U}}$ since $1+\eta_{N}^{0}\geq 1/2>0$, and $\left|\bm{U}_{N}^{0}\right|_{H^{s}}\approx 1$. Notice also that $\bm{U}_{N}^{0}\notin{\mathcal{U}}_{{\mathscr{H}}}$ since $1+\eta_{N}^{0}(\frac{\pi}{2p})-|u_{N}^{0}(\frac{\pi}{2p})|^{2}=0$, but one could enforce $\bm{U}_{N}^{0}\in{\mathcal{U}}_{{\mathscr{H}}}$ while keeping valid all previous statements by considering e.g. $u_{p}(x)=\frac{1}{2}\sin(px)$. In that case, the Hamiltonian structure allows to propagate the functional ${\mathcal{F}}_{{\mathscr{H}},s}(\bm{U})\approx\big\lvert\bm{U}_{N}\big\rvert_{H^{s}}$ with